Circular No. 8904

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 Y1 (LINEAR) An apparently asteroidal object discovered by the LINEAR survey (dis-

covery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been reported to show cometary appearance by several CCD astrometrists. K. Sárneczky writes that a 960-s co-added unfiltered image taken with the 0.60-m Schmidt telescope at Konkoly Observatory on Dec. 17.15 UT shows a 10′′ coma and a 15′′-long tail in p.a. 220^{o}. E. Guido and G. Sostero (Castellammare di Stabia, Italy) report that thirty co-added 60-s unfiltered exposures obtained remotely using a 0.25-m f/3.4 reflector near Mayhill, NM, U.S.A., on Dec. 17.41–17.47 show a compact coma nearly 8′′ in diameter, slightly elongated toward the southwest; a similar co-adding of exposures from Dec. 18.46–18.47 shows the presence of a coma nearly 10′′ in diameter. J. G. Ries writes that, in images obtained in less-than-ideal conditions with the McDonald Observatory 0.76-m reflector (+ prime-focus corrector) on Dec. 18.45–18.47, the object seems extended, showing a slight asymmetry toward the southwest.

2007 UT  α_{2000}  δ_{2000}  Mag.
Dec. 16.37807  10^{h}04^{m}42.^{s}65  −15^{o}05′40′′.6  18.8

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-Y22.

T = 2008 Mar. 10.736 TT  ω = 354.^{o}944
Ω = 133.191 }2000.0
q = 3.38036 AU  i = 110.137

V458 VULPECULAE T. R. Prater, R. J. Rudy, D. K. Lynch, T. R. Prater, and S. Mazuk, The

Aerospace Corporation; R. B. Perry, Langley Research Center, NASA; and R. C. Puetter, University of California, San Diego, report on 0.47- to 2.5-μm spectroscopy of V458 Vul (cf. IAUC 8861, 8863, 8883) that was obtained on Dec. 15 UT using VNIRIS on the Lick 3-m telescope. The object’s emission spectrum has moved rapidly to higher excitation and shows weak unidentified novae lines and strong He I and II lines. Also present were [Fe VII] lines at 572.1 and 608.7 nm. The spectrum showed a variety of line profiles: Gaussian, flat-topped, and doubled.

2007 December 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green